OpenStudy (anonymous):

Factor the following completely Im not sure how to do this please show work I will give medals 1. 3x^2 + 8x – 16 2. c^2 – 121 3. p^2 + 22p + 40

OpenStudy (luigi0210):

Well 2 is the difference of two perfect squares.

OpenStudy (anonymous):

3*-16=-48 12-4=8 12*-4=-48

OpenStudy (anonymous):

is that the answer for the first one?

OpenStudy (anonymous):

yes,write 8x=12x-4x and make factors.

OpenStudy (anonymous):

im confused

OpenStudy (whpalmer4):

Split the middle term: \[3x^2 + 8x -16 = 3x^2 + 12x - 4x -16\]Now rewrite as two groups: \[(3x^2+12x) - (4x+16)\](notice the sign change on the 16 as I factored out the minus sign!) Now factor each group separately: \[3x(x+4) - 4(x+4)\]Notice that each of those expressions has a common factor? \[(x+4)(3x-4)\] Check the work: \[(x+4)(3x-4) = x(3x-4) + 4(3x-4) = 3x^2 - 4x + 12x -16 \]\[\qquad= 3x^2 +8x -16\checkmark\]

OpenStudy (anonymous):

thanks. what about the other two?

OpenStudy (whpalmer4):

You come up with the coefficients to split it by multiplying the coefficient of the 1st term with the coefficient of the 3rd term: 3*-16 = -48. Now find a pair of factors of -48 that add to 8: 12 and -4 are that pair. You can use the same procedure on the last one: \[p^2 + 22p + 40\]1*40 = 40. Factors of 40 that add to 22 are 2 and 20. I'll let you do the rest and check your work. @Luigi0210 already told you how to do the 2nd one. \[a^2-b^2 = (a+b)(a-b)\]